Consider the conic represented by 4(x²+2y)=9. Now shift the origin to (1,-1) and then rotate the axes through π/3 What is the resultant equation? What geometrical object does it represent?

1
Expert's answer
2022-06-30T12:31:13-0400

"4(x^2+2y)=9\implies y=\frac{9}{8}-\frac{1}{2}x^2" , hence the conic is a parabola.

Let "(x_1,y_1)"  be the new coordinates of point "(x,y)"  after the shift of the origin and let "(x_2,y_2)"  be the new coordinates of "(x_1,y_1)"  after the rotation, then

"x=x_1+1"

"y=y_1-1"

"x_1=x_2cos(\pi/3)-y_2sin(\pi/3)=\frac{1}{2}x_2-\frac{\sqrt{3}}{2}y_2"

"y_1=x_2sin(\pi/3)+y_2cos(\pi/3)=\frac{\sqrt{3}}{2}x_2+\frac{1}{2}y_2"

hence

"x=\frac{1}{2}x_2-\frac{\sqrt{3}}{2}y_2+1"

"y=\frac{\sqrt{3}}{2}x_2+\frac{1}{2}y_2-1"

substituting the last two equations into the original equation:

"4((\frac{1}{2}x_2-\frac{\sqrt{3}}{2}y_2+1)^2+2(\frac{\sqrt{3}}{2}x_2+\frac{1}{2}y_2-1))=9,"

"(x_2-\sqrt{3}y_2+2)^2+4\sqrt{3}x_2+4y_2-8=9,"

"x_2^2+3y_2^2+4-2\sqrt{3}x_2y_2+4x_2-4\sqrt{3}y_2+\\
4\sqrt{3}x_2+4y_2-8=9,"

"x_2^2+3y_2^2-2\sqrt{3}x_2y_2+4(1+\sqrt{3})x_2+\\4(1-\sqrt{3})y_2=13."